# Boundary controllability of the Korteweg-de Vries equation: The Neumann case

### v2

@@ Abstract
This article gives a necessary first step to understanding the critical set phenomenon for the Korteweg-de Vries (KdV) equation posed on interval $[0,L]$ considering the Neumann boundary conditions with only one control input. We showed that the KdV equation is controllable in the critical case, i.e., when the spatial domain $L$ belongs to the set $\mathcal{R}_{c}$, where $c\neq-1$ and
@@ 1 Introduction
We had known when we formulated the waves as a free boundary problem of the incompressible, irrotational Euler equation in an appropriate non-dimensional form, there exist two non-dimensional parameters $\delta:=\frac{h}{\lambda}$ and $\varepsilon:=\frac{a}{h}$, where the water depth, the wavelength and the amplitude of the free surface are parameterized as $h,\lambda$ and $a$, respectively. See, for instance, [2, 3, 1, 4, 23, 26] and references therein for a rigorous justification. Moreover, another non-dimensional parameter $\mu$ appears, the Bond number, to measure the importance of gravitational forces compared to surface tension forces also appears. Considering the physical condition $\delta\ll 1$ we can characterize the waves, called long waves or shallow water waves. In particular, considering the relations between $\varepsilon$ and $\delta$, we can have the following regime:
@@
    ±$\pm 2\eta_{t}+3\eta\eta_{x}+\left(\frac{1}{3}-\mu\right)\eta_{xxx}=0.$
Today, it is well known that this equation has an important phenomenon that directly affects the control problem related to them, the so-called critical length phenomenon. Let us briefly present the control problem, which makes the phenomenon of critical lengths emerge. The control problem was presented in a pioneering work of Rosier [24] that studied the following system
$\left\{\begin{array}[c]{lll}u_{t}+u_{x}+uu_{x}+u_{xxx}=0,&&\text{ in }(0,L)\times(0,T),\\ u(0,t)=0,\text{ }u(L,t)=0,\text{ }u_{x}(L,t)=g(t),&&\text{ in }(0,T),\\ u(x,0)=u_{0}(x),&&\text{ in }(0,L),\end{array}\right.$ (1.1)
@@
$\left\{\begin{array}[c]{lll}u_{t}+u_{x}+u_{xxx}=0,&&\text{ in }(0,L)\times(0,T),\\ u(0,t)=0,\text{ }u(L,t)=0,\text{ }u_{x}(L,t)=g(t),&&\text{ in }(0,T),\\ u(x,0)=u_{0}(x)&&\text{ in }(0,L),\end{array}\right.$ (1.3)
is controllable; roughly speaking, if $L\in\mathcal{N}$ system (1.3) is not controllable, that is, there exists a finite-dimensional subspace of $L^{2}(0,L)$, denoted by $\mathcal{M}=\mathcal{M}(L)$, which is unreachable from 00 for the linear system. More precisely, for every nonzero state $\psi\in\mathcal{M}$, $g\in L^{2}(0,T)$ and $u\in C([0,T];L^{2}(0,L))\cap L^{2}(0,T;H^{1}(0,L))$ satisfying (1.3) and $u(\cdot,0)=0$, one has $u(\cdot,T)\neq\psi$.
@@ 1.1 Problem set
$\left\{\begin{array}[c]{lll}u_{t}+u_{x}+u_{xxx}+uu_{x}=0,&&\text{ in }(0,L)\times(0,T),\\ u_{xx}(0,t)=u_{xx}(L,t)=0,&&\text{ in }(0,T),\\ u_{x}(L,t)=h(t),&&\text{ in }(0,T),\\ u(x,0)=u_{0}(x),&&\text{ in }(0,L),\end{array}\right.$ (1.4)
where $h(t)$ will be considered as a control input. Recently, the first author dealt with the control problem related to the system (1.4). Precisely, was proved that their solutions are exactly controllable in a neighborhood of $c$ if the length $L$ of the spatial domain $(0,L)$ does not belong to the set
@@
$\left\{\begin{array}[c]{lll}v_{t}+(1+c)v_{x}+v_{xxx}=0,&\text{ \ in \ }(0,L)\times(0,T),\\ v_{xx}(0,t)=v_{xx}(L,t)=0,&\text{ \ in \ }(0,T),\\ v_{x}(L,t)=h(t),&\text{ \ in \ }(0,T),\\ v(x,0)=v_{0}(x),&\text{ \ in \ }(0,L).\end{array}\right.$ (1.5)
Precisely the authors in [7] proved the following result.
@@
For $c\neq-1$, the linear system (1.5) is exactly controllable in the space $L^{2}(0,L)$ if and only if $L\notin\mathcal{R}_{c}$. Otherwise, that is, if $c=-1$, the system (1.5) is not exactly controllable in the space $L^{2}(0,L)$ for any $L>0$.
With this in hand, they extend the result for the nonlinear system (1.4) using fixed point argument achieving Theorem 1.2 whenever $L\notin\mathcal{R}_{c}$. It is important to point out that fixing $k\in\mathbb{N}^{*}$ and considering $m=l=k$ we have $L=2k\pi$, when $c=0$, and, from Theorem 1.3, it follows that (1.5) are not exactly controllable. Additionally, as before mentioned, we do not know if the system (1.4) is exactly controllable. So, in this context, the natural questions appear:
Question $\mathcal{B}$: Given $T>0$, $L\in\mathcal{R}_{c}$ and $y_{0},y_{T}\in L^{2}(0,L)$ close enough to $c$, can one find an appropriate control input $h\in L^{2}(0,T)$ such that the solution $y$ of the system (1.9), corresponding to $h$ and $y_{0}$, satisfies $y(\cdot,T)=y_{T}$?
Question $\mathcal{C}$: Given $T>0$, $L\in\mathcal{R}_{c}$ and $y_{0},y_{T}\in L^{2}(0,L)$. The system (1.9) is exactly controllable in the critical length $L$, that is when $L\in\mathcal{R}_{c}$?
@@ 1.2 Main results
Let us consider the following nonlinear control system
$\left\{\begin{array}[]{lll}y_{t}+(1+c)y_{x}+y_{xxx}+yy_{x}=0,&\text{ in }(0,L)\times(0,T),\\ y_{xx}(0,t)=0,\ y_{x}(L,t)=h(t),\ y_{xx}(L,t)=0,&\text{ in }(0,T),\\ y(x,0)=y_{0}(x),&\text{ in }(0,L),\end{array}\right.$ (1.9)
where $h$ will be as a control input and $y$ is the state. Theorem 1.2 says that when $L\notin\mathcal{R}_{c}$, system (1.9) is locally controllable around $c$ but we do not know if the same holds when $L\in\mathcal{R}_{c}$. The main result in this work provides an affirmative answer to the Question $\mathcal{C}$. Precisely, we have the following:
@@
Let $T>0$, $c=0$ and $L\in\mathcal{R}_{0}$. Then, system (1.9) is exactly controllable around the origin 00 in $L^{2}(0,L)$, that is, there exists $\delta>0$ such that, for very $y_{0},y_{T}\in L^{2}(0,L)$ with
$\|y_{0}\|_{L^{2}(0,L)},\|y_{T}\|_{L^{2}(0,L)}<\delta$
it is possible to find $h\in L^{2}(0,T)$ such that the corresponding solution of (1.9) satisfying $y(\cdot,0)=y_{0}$ and $y(\cdot,T)=y_{T}$.
@@
To prove the previous results we need an auxiliary property that ensures that for $c$ near enough to 00 (small perturbations of 00), the system (1.4) is exactly controllable in a neighborhood of $c$ in $\mathcal{R}_{c})$, or precisely, for $d$ close enough to $c$ one has $L\notin\mathcal{R}_{d}$ so that, the system (1.5) corresponding to $d$ is exactly controllable, answering the Question $\mathcal{B}$. In other words, the set of critical lengths is sensitive to small disturbances in equilibrium $c$, and the result can be read as follows.
@@
Let $T>0$, $c\neq-1$ and $L\in\mathcal{R}_{c}$. There exists $\varepsilon_{c}>0$ such that, for every $d\in(c-\varepsilon_{c},c+\varepsilon_{c})\backslash\{c\}$, $d\neq-1$, we have $L\notin\mathcal{R}_{d}$. Consequently, the linear system (1.5), with $c=d$, is exactly controllable; and the nonlinear system (1.9) is exactly controllable around the steady state $d$ in $L^{2}(0,L)$, that is, there exists $\delta_{d}>0$ such that, for any $y_{0},y_{T}\in L^{2}(0,L)$ with
$\|y_{0}-d\|_{L^{2}(0,L)}<\delta_{d}\text{ \ and \ }\|y_{T}-d\|_{L^{2}(0,L)}<\delta_{d},$
one can find $h\in L^{2}(0,T)$ such that the system (1.9) admits a unique solution $y\in\mathcal{Z}_{T}$ satisfying $y(\cdot,0)=y_{0}$ and $y(\cdot,T)=y_{T}$.
@@ 1.3 Heuristic and paper’s outline
The proof of the Theorem 1.6 is based on the topological properties of real numbers together with the Theorem 1.2. Moreover, with this in hand, both results stated in the previous paragraph (Theorems 1.4 and 1.5) rely on the so-called return method together with the fixed point argument.
It is important to point out that the return method was introduced by J.M. Coron in [12] (see also [13]) and has been used by several authors to prove control results in the critical lengths for the KdV-type equation (see, for instance, [8, 10, 11, 15]). This method consists of building particular trajectories of the system (1.9) starting and ending at some equilibrium such that the linearization of the system around these trajectories has good properties. Here, we use a combination of this method with a fixed point argument, successfully applied in [16]. We mention that this method can be applied together with quasi-static deformations and power series expansion, we infer to the reader the nice book of Coron [14] about more details of the method.
Concerning the construction of solutions to the Theorems 1.4 and 1.5 we follow the following procedure: In the first time, we construct a solution that starts from $y_{0}$ and reaches at time $T/3$ a state which is in some sense close to $d$ (which is yet to be defined). Then we construct a solution (close to the state solution $d$ ), which starts at time $32T/$ 3 from the previous state. In the last step, we bring the latter state to 0 via a function $y_{2}$, as we can see in Figure 1 below. For details of this construction see the characterization of the function $y$ in (3.52).
@@
We finish this introduction with an outline of this work which consists of four parts, including the introduction. Section 2 gives an overview of the well-posedness of the system (1.9). Section 3 is devoted to proving carefully the controllability of the system (1.9) when $L\in\mathcal{R}_{c}$. Precisely, in the first part of Section 3, we deal with the proof of Theorem 1.6. In the second part, we prove the construction of the function $y$ mentioned before, and finally, in the third part of Section 3, we use these previous results to achieve Theorem 1.4.
@@ 2 Overview of the well-posedness theory
[…]
@@
Using the Proposition 2.1 we can get properties for the linear problem
@@
[…]
@@
With these previous results in hand, the following proposition gives us the well-posedness to the general system (2.17), which will be used several times so, for the sake of completeness, we will give the proof.
@@
Thus, $\Gamma:B\rightarrow B$ is a contraction so that, by Banach’s fixed point theorem, $\Gamma$ has a fixed point $y\in B$ which is a solution to the problem (2.17) in $[0,\theta]$, corresponding to data $(y_{0},{\bf h},f)$. Additionally, inequalities (2.18) and (2.19) yields that
@@
Hence, $\Gamma:B\rightarrow B$ is a contraction so that, by Banach’s fixed point theorem, $\Gamma$ has a fixed point $y\in B$ which is a solution to the problem (2.24) in $[0,\theta]$, corresponding to data $(y_{0},{\bf h},f)$.
@@
respectively. Then, $w=y-u$ solves the problem
@@
Hence, the constant $D$ can be chosen depending only on $T$ and $\|a\|_{\mathcal{Z}_{T}}$ (and also on $\lambda$), so that, (2.28) gives us the uniqueness of the solution which turns the solution map $S$ well-defined. Moreover, from (2.28)
@@
    For every $a\in\mathcal{Z}_{T}$, the Proposition 2.4 give us the well-definition for the solution operator
    $\Lambda_{a}:L^{2}(0,L)\times\mathcal{H}_{T}\times L^{1}(0,T;L^{2}(0,L))\rightarrow\mathcal{Z}_{T}$
    where, for each $(y_{0},{\bf h},f)\in L^{2}(0,L)\times\mathcal{H}_{T}\times L^{1}(0,T;L^{2}(0,L))$, $\Lambda_{a}(y_{0},{\bf h},f)$ is the corresponding solution to the problem (2.17). Furthermore, the Proposition 2.6 guarantees that $\Lambda_{a}$ is a bounded linear operator.
@@
3 Boundary controllability in the critical lenght
@@
when $L$ is a critical length. We will use the return method together with the fixed point argument to ensure the controllability of the system (3.32) when $L\in\mathcal{R}_{c}$. Before presenting the proof of the main result, let us give a preliminary result which is important in our analysis.
@@ 3.1 An auxiliary result
is not exactly controllable. So, in this section, we first give the proof of Theorem 1.6.
@@
From all the above we conclude that
@@
Therefore for $d\in(c-\epsilon_{1},c+\epsilon_{1})\backslash\{c\}$ we have that $d\notin\mathcal{A}_{1}\cup\mathcal{B}_{1}$ and, therefore, $L\notin\mathcal{R}_{d}$.
@@
Proceeding as in the first case, we conclude that there exists $\epsilon_{2}>0$ such that, for every $d\in(c-\epsilon_{2},c+\epsilon_{2})\backslash\{c\}$ we have $d\notin\mathcal{A}_{2}\cup\mathcal{B}_{2}$ so that $L\notin\mathcal{R}_{d}$. Considering $\epsilon_{c}:=\min\{\epsilon_{1},\epsilon_{2}\}$, thanks thanks to the [7, Proposition 3.6] and Theorem 1.2, the proof is completed.

### v3

@@ Abstract
This article gives a necessary first step to understanding the critical set phenomenon for the Korteweg-de Vries (KdV) equation posed on the interval $[0,L]$, considering the Neumann boundary conditions with only one control input. We showed that the KdV equation is controllable in the critical case, i.e., when the spatial domain $L$ belongs to the set $\mathcal{R}_{c}$, where $c\neq-1$ and
@@ 1 Introduction
We had known when we formulated the waves as a free boundary problem of the incompressible, irrotational Euler equation in an appropriate non-dimensional form, there exist two non-dimensional parameters $\delta:=\frac{h}{\lambda}$ and $\varepsilon:=\frac{a}{h}$, where the water depth, the wavelength and the amplitude of the free surface are parameterized as $h,\lambda$ and $a$, respectively. See, for instance, [2, 3, 1, 4, 23, 26] and references therein for a rigorous justification. Moreover, another non-dimensional parameter $\mu$ appears, the Bond number, to measure the importance of gravitational forces compared to surface tension forces. Considering the physical condition $\delta\ll 1$, we can characterize the waves, called long waves or shallow water waves. In particular, considering the relations between $\varepsilon$ and $\delta$, we can have the following regime:
@@
    $\pm 2\eta_{t}+3\eta\eta_{x}+\left(\frac{1}{3}-\mu\right)\eta_{xxx}=0.$
Today, it is well known that this equation has an important phenomenon that directly affects the control problem related to it, the so-called critical length phenomenon. Let us briefly present the control problem, which makes the phenomenon of critical lengths emerge. The control problem was presented in a pioneering work of Rosier [24] that studied the following system
[…]
@@
[…]
is controllable; roughly speaking, if $L\in\mathcal{N}$ system (1.3) is not controllable, that is, there exists a finite-dimensional subspace of $L^{2}(0,L)$, denoted by $\mathcal{M}=\mathcal{M}(L)$, which is unreachable from 0 for the linear system. More precisely, for every nonzero state $\psi\in\mathcal{M}$, $g\in L^{2}(0,T)$ and $u\in C([0,T];L^{2}(0,L))\cap L^{2}(0,T;H^{1}(0,L))$ satisfying (1.3) and $u(\cdot,0)=0$, one has $u(\cdot,T)\neq\psi$.
@@ 1.1 Problem set
[…]
where $h(t)$ will be considered as a control input. Recently, the first author dealt with the control problem related to the system (1.4). Precisely, it was proved that their solutions are exactly controllable in a neighborhood of $c$ if the length $L$ of the spatial domain $(0,L)$ does not belong to the set
@@
[…]
Precisely, the authors in [7] proved the following result.
@@
For $c\neq-1$, the linear system (1.5) is exactly controllable in the space $L^{2}(0,L)$ if and only if $L\notin\mathcal{R}_{c}$. Otherwise, that is, if c=1c=-1, the system (1.5) is not exactly controllable in the space $L^{2}(0,L)$ for any $L>0$.
With this in hand, they extend the result for the nonlinear system (1.4) using a fixed point argument, achieving Theorem 1.2 whenever $L\notin\mathcal{R}_{c}$. It is important to point out that fixing $k\in\mathbb{N}^{*}$ and considering $m=l=k$, we have $L=2k\pi$, when $c=0$, and, from Theorem 1.3, it follows that (1.5) are not exactly controllable. Additionally, as mentioned before, we do not know if the system
$\left\{\begin{array}[]{lll}u_{t}+(1+c)u_{x}+u_{xxx}+uu_{x}=0,&\text{ in }(0,L)\times(0,T),\\ u_{xx}(0,t)=0,\ u_{x}(L,t)=h(t),\ u_{xx}(L,t)=0,&\text{ in }(0,T),\\ u(x,0)=u_{0}(x),&\text{ in }(0,L),\end{array}\right$. (1.9)
is exactly controllable. So, in this context, the natural questions appear:
Question $\mathcal{B}$: Given $T>0$, $L\in\mathcal{R}_{c}$ and $u_{0},u_{T}\in L^{2}(0,L)$ close enough to $c$, can one find an appropriate control input $h\in L^{2}(0,T)$ such that the solution $u$ of the system (1.9), corresponding to $h$ and $u_{0}$, satisfies $u(\cdot,T)=u_{T}$?
Question $\mathcal{C}$: Given $T>0$, $L\in\mathcal{R}_{c}$ and $u_{0},u_{T}\in L^{2}(0,L)$. The system (1.9) is exactly controllable in the critical length $L$, that is, when $L\in\mathcal{R}_{c}$?
@@ 1.2 Main results
Let us consider the nonlinear control system (1.9) with $h$ as a control input and $u$ as the state. Theorem 1.2 says that when $L\notin\mathcal{R}_{c}$, system (1.9) is locally controllable around $c$, but we do not know if the same holds when $L\in\mathcal{R}_{c}$. The main result in this work provides an affirmative answer to the Question $\mathcal{C}$. Precisely, we have the following:
@@
Let $T>0$, $c=0$ and $L\in\mathcal{R}_{0}$. Then, system (1.9) is exactly controllable around the origin 0 in $L^{2}(0,L)$, that is, there exists $\delta>0$ such that, for very $u_{0},u_{T}\in L^{2}(0,L)$ with
$\|u_{0}\|_{L^{2}(0,L)},\|u_{T}\|_{L^{2}(0,L)}<\delta,$
it is possible to find $h\in L^{2}(0,T)$ such that the corresponding solution of (1.9) satisfying $u(\cdot,0)=u_{0}$ and $u(\cdot,T)=u_{T}$.
@@
To prove the previous results we need an auxiliary property that ensures that for $c$ near enough to 0 (small perturbations of 0), the system (1.4) is exactly controllable in a neighborhood of $c$ in $\mathcal{R}_{c})$, or precisely, for $d$ close enough to $c$ one has $L\notin\mathcal{R}_{d}$ so that, the system (1.5) corresponding to $d$ is exactly controllable, answering the Question $\mathcal{B}$. In other words, the set of critical lengths is sensitive to small disturbances in equilibrium $c$, and the result can be read as follows.
@@
[…]
$\|u_{0}-d\|_{L^{2}(0,L)}<\delta_{d}\text{ \ and \ }\|u_{T}-d\|_{L^{2}(0,L)}<\delta_{d},$
one can find $h\in L^{2}(0,T)$ such that the system (1.9) admits a unique solution $u\in\mathcal{Z}_{T}$ satisfying $u(\cdot,0)=u_{0}$ and $u(\cdot,T)=u_{T}$.
@@ 1.3 Heuristic and paper’s outline
The proof of Theorem 1.6 is based on the topological properties of real numbers together with Theorem 1.2. Moreover, with this in hand, both results stated in the previous paragraph (Theorems 1.4 and 1.5) rely on the so-called return method together with the fixed point argument.
It is important to point out that the return method was introduced by J.M. Coron in [12] (see also [13]) and has been used by several authors to prove control results in the critical lengths for the KdV-type equation (see, for instance, [8, 10, 11, 15]). This method consists of building particular trajectories of the system (1.9) starting and ending at some equilibrium such that the linearization of the system around these trajectories has good properties. Here, we use a combination of this method with a fixed point argument, successfully applied in [16]. We mention that this method can be applied together with quasi-static deformations and power series expansion. We refer the reader to the nice book of Coron [14] for more details of the method.
It is important to emphasize that our control $h\in L^{2}(0,T)$ is guaranteed by the construction of $u$ (see below). In this setting, the resulting control (or trajectory) is not unique. Furthermore, choosing smoother initial and terminal data yields higher regularity for both the control and the state within this framework. However, as is well known in the literature, when the initial data lie in $L^{2}$, the regularity of the control is optimal at the $L^{2}$–level.
Concerning the construction of solutions to the Theorems 1.4 and 1.5, we follow the following procedure: In the first step, we construct a solution that starts from $y_{0}$ and reaches at time $T/3$ a state which is in some sense close to $d$ (which is yet to be defined). Then we construct a solution (close to the state solution $d$ ), which starts at time $2T/3$ from the previous state. In the last step, we bring the latter state to 0 via a function $u_{2}$, as we can see in Figure 1 below, considering the characterization of the function $u:[0,L]\times[0,T]\rightarrow\mathbb{R}$ by
$u=\left\{\begin{array}[]{ll}u_{1},&\text{ in }[0,T/3],\\ d,&\text{ in }[T/3,2T/3],\\ u_{2},&\text{ in }[2T/3,T].\end{array}\right$.
@@
We finish this introduction with an outline of this work, which consists of three parts, including the introduction. Section 2 gives an overview of the well-posedness of the system (1.9). Section 3 is devoted to proving carefully the controllability of the system (1.9) when $L\in\mathcal{R}_{c}$. Precisely, in the first part of Section 3, we deal with the proof of Theorem 1.6. In the second part, we prove the construction of the function $u$ mentioned before, and, in the third part of Section 3, we use these previous results to achieve Theorem 1.4. Finally, Section 4 is devoted to final remarks and comments.
@@ 2 Overview of the well-posedness theory
${\bf h}=0\text{ in }(\theta,T]\quad\text{and}\quad f=0\text{ in }(\theta,T].$
@@
Using the Proposition 2.1, we can get properties for the linear problem
@@
$\int_{0}^{T}\|uv_{x}(\cdot,t)\|_{L^{2}(0,L)}dt\leq C\left(T^{\frac{1}{2}}+T^{\frac{1}{3}}\right)\|u\|_{\mathcal{Z}_{T}}\|v\|_{\mathcal{Z}_{T}},$
@@
With these previous results in hand, the following proposition gives us the well-posedness of the general system (2.17), which will be used several times, so, for the sake of completeness, we will give the proof.
@@
Thus, $\Gamma:B\rightarrow B$ is a contraction so that, by Banach’s fixed point theorem, $\Gamma$ has a fixed point $y\in B$ which is a solution to the problem (2.17) in $0,\theta]$, corresponding to data $(y_{0},{\bf h},f)$. Additionally, inequalities (2.18) and (2.19) yields that
@@
Hence, $\Gamma:B\rightarrow B$ is a contraction so that, by Banach’s fixed point theorem, $\Gamma$ has a fixed point $y\in B$ which is a solution to the problem (2.24) in $0,\theta]$, corresponding to data $(y_{0},{\bf h},f)$.
@@
respectively. Then, w=yuw=y-u solves the problem
@@
Hence, the constant $D$ can be chosen depending only on $T$ and $\|a\|_{\mathcal{Z}_{T}}$ (and also on $\lambda$), so that (2.28) gives us the uniqueness of the solution, which turns the solution map $S$ well-defined. Moreover, from (2.28)
@@
    For every $a\in\mathcal{Z}_{T}$, the Proposition 2.4 give us the well-definition for the solution operator Λ_a:L^2(0,L)×H_T×L^1(0,T;L^2(0,L))Z_T where, for each $(y_{0},{\bf h},f)\in L^{2}(0,L)\times\mathcal{H}_{T}\times L^{1}(0,T;L^{2}(0,L))$, $\Lambda_{a}(y_{0},{\bf h},f)$ is the corresponding solution to the problem (2.17). Furthermore, the Proposition 2.6 guarantees that $\Lambda_{a}$ is a bounded linear operator.
@@
3 Boundary controllability in the critical length
@@
when $L$ is a critical length. We will use the return method together with the fixed point argument to ensure the controllability of the system (3.32) when $L\in\mathcal{R}_{c}$. Before presenting the proof of the main result, let us give a preliminary result that is important in our analysis.
@@ 3.1 An auxiliary result
it is not exactly controllable. So, in this section, we first give the proof of Theorem 1.6.
@@
From all the above, we conclude that
@@
[…]
@@
[…]
Remark 3.2.
It can seen that, for $c\neq-1$ and $L\in\mathcal{R}_{c}$, we must have $0<\epsilon_{c}\leq\frac{\pi^{2}}{3L^{2}}$. Hence, $\epsilon_{c}$ does not depend on the control time $T$ and $\epsilon_{c}\rightarrow 0$ when $L\rightarrow\infty$. On the other hand, for $L$ close enough to 0, the tolerance $\epsilon_{c}$ can be considered extremely large.